I need help with all three problems

EXPERIMENT # 2: ORGANIC STRUCTURES AND MOLECULAR MODELS B. Bond Lengths and Resonance TeGaussian molecular modeling software utilizes the sophisticated Molecular Orbital Theory to predict the optimal geometry of molecules in the gas phase. Techniques used are theoretical predictions rather than experimental results. With the help of your instructor, use the Gaussian program to investigate the bond lengths of ethane (C,He), ethylene (CH), acetylene (C H2), and benzene (C&Ha). In each case build the molecule, optimize it in the calculations menu, and then measure all carbon-carbon bond lengths. Draw the molecules here and label the bond lengths. i. ii. Describe the relationship between the type of bond (single, double, triple) and the length of the bonds based on the results of your calculations. Which bond is predicted to be the strongest bond? ii. How can the concept of resonance from Lewis Theory be reconciled with the ictions you found for benzene using Molecular Orbital Theory?
Answer #1

ethane : C-C = 1.520 Angstroms(A) single bond because longest C-C bonds C-H = 1.110 A

ethene : C-C = 1.334 A double bond because intermediate C-C bond length C-H = 1.085 A

ethyne : C-C = C-C = 1.20 A triple bond because shortes C-C bond length C-H = 1.058 A

benzene : C-C = 1.399 partial double bond, bond length lies between that of a double and single bond length which expalins resonance as explained in lewis theory. (part3 of question) C-H = 1.087 A

H6 H3 H2 H1H3 C2 H2 H4

C-H bond length also changes as hybridisation changes. longest C-H bond is for sp3 hybridised carbon in ethane and shortest in sp hybridised carbon in ethyne.